What is the p-value for a two tailed test when the test statistic is z=0.7?

0.484

0.089

0.242

0.944
Answer #1

solution

P(z >0.7) = 1 - P(z < 0.7 ) = 1-0.758=0.242 z table

P-value = 2*0.242=0.4840